Cost Flow Relationships

The following information is available for the first month of operations of Bahadir Company, a manufacturer of mechanical pencils:

Sales $398,990
Gross profit 232,610
Cost of goods manufactured 199,500
Indirect labor 86,580
Factory depreciation 13,170
Materials purchased 122,890
Total manufacturing costs for the period 229,420
Materials inventory, ending 16,360

Using the above information, determine the following missing amounts:

a. Cost of goods sold $
b. Finished goods inventory at the end of the month $
c. Direct materials cost $
d. Direct labor cost $
e. Work in process inventory at the end of the month $
Answer #1

Sales - Cost of Goods Sold = Gross Profit

Cost of Goods Sold = Sales - Gross profit

= 398,990 - 232,610

= 166,380

Finished goods inventory = Cost of Goods manufactured - Cost of Goods Sold

= 199,500 - 166,380

= 33,120

Direct materials cost - Materials purchased - Materials inventory, ending

= 122,890 - 16,360

= 106,530

Total manufacturing cost = Direct Materials + Direct Labor + Overhead

Direct labor = Total manufacturing cost - Direct materials - Overhead

= 229,420 - 106,530 - (indirect labor 86,580+factory depreciation 13,170)

= 23,140

Work in Process Inventory at the end of month

= Total manufacturing cost - Cost of Goods manufactured

= 229,420 - 199,500

= 29,920